Verify Green’s theorem in the plane for $ (xy + Sy?)dx + 3xdy,

where c is the closed curve of the region bounded by y = x and y = x?.


1
Expert's answer
2021-05-26T14:01:50-0400

Green’s theorem:

"\oint_CPdx+Qdy=\iint_D(Q_x-P_y)dA"


We have:

"P=xy+y,Q=3x"

Then:

"Q_x=3,P_y=x+1"

"\oint_C(xy+y)dx+3xdy=\int^1_0(2-x)dx\int^{1-x}_xdy=\int^1_0(2-x)(1-2x)dx="

"=(2x-5x^2\/2+2x^3\/3)|^1_0=2-5\/2+2\/3=1\/6"




"\oint_C(xy+y)dx+3xdy=\int_1^{0.5}[(x(1-x)+1-x)dx+3xd(1-x)]+"

"+\int_{0.5}^{0}[(x^2+x)dx+3xdx]=\intop^{0.5}_1(1-x^2-3x)dx+\int^0_{0.5}(x^2+4x)dx="

"=(x-x^3\/3-3x^2\/2)|^{0.5}_1+(x^3\/3+2x^2)|^0_{0.5}="


"=0.25-0.125\/3-0.75\/2-1+1\/3+1.5-0.125\/3-0.5=1\/4-1\/12-3\/8+1\/3=1\/6"


We have the same results, so Green’s theorem is verified.